# thanks: These two authors contributed equally to this workthanks: These two authors contributed equally to this work

### v1

@@ VI.2 The Fitting of Critical Exponents
where $b\cdot L^{-\omega}$ refers to the finite-size correction and $c$ derives from the analytic part of freedom energy. During the fitting process, various values of $\omega$ are tested. It is observed that when $\omega$ is excessively large or small, the data cannot be adequately fitted. However, when $\omega$ falls within a certain range, an appropriate fit yields reliable estimates for exponents. The details and results of the fitting are shown in Tables 6, 7 and 5, respectively. Figure 15 and 16 visually show the reliability of the fitting.
Table 5: Estimates of two exponents $y_{t}$ and $\eta$ for non-classical regime. The results are obtained from the fitting: Eq. (41) and (42).
$\sigma$       $y_{t}$       $\eta$
@@
2 0.37(4) 0.260(5)
To be specific, when taking the estimated value of $\eta$ and $y_{t}$, according to the fitting, $\chi L^{\eta-2}-bL^{-\omega}$ and $KL^{-y_{t}}-bL^{-\omega}$ should converge to a non-zero value as $L$ increases; otherwise they’ll either diverge or reduce to 0. It can be seen that in both figures, curves corresponding to the estimated value of exponents finally converge to a plateau, indicating the reliability of the fitting results. Figure 17 presents our numerical estimates of critical exponents at different $\sigma$. Panel (a) displays our estimates for $\eta$ across different $\sigma$ values, alongside the estimates by Picco[11] and Luijten[8].
Figure 17: The value of critical exponent at various $\sigma$. (a) The red dots denote the estimates of $\eta$ in this work. Picco and Luijten’s results are also contained [11, 8], respectively denoted by blue and green dots. Although the model they studied is the Ising model, our estimated values of $\eta$ are consistent with their results. The dotted lines correspond to Sak’s prediction: $\eta=\rm{max}(2-\sigma,\frac{1}{4})$. Obviously, our results deviate from Sak’s prediction, suggesting that for $21.75<\sigma<$ 2, the short-range behavior is not recovered. (b) The red dots denote the estimates of $y_{t}$ in this work. The dotted line represents the results from second-order $\epsilon$ expansion [2] (note that $y_{t}=1/\nu$).
@@
Despite the difference in models (Luijten and Picco studied the 2D LR Ising model, whereas we are studying the 2D LR XY model), our results are consistent with Picco’s, which does not agree with Sak’s prediction. In our results, $\eta$ interpolates smoothly between $\eta=2-\sigma$ for $\sigma$ near 1 and $\eta=\eta_{SR}=\frac{1}{4}$ for $\sigma>2$. For $\sigma<2$, $\eta$ remains significantly higher than $\frac{1}{4}$, indicating the phase transition doesn’t belong to the short-range university class. For $\sigma=2$, the estimated value of $\eta$ still slightly exceeds $\frac{1}{4}$, which could be attributed to some logarithmic corrections in the behavior of $\chi$ at the marginal point $\sigma=2$. As mentioned before, Ref. [9] suggests that Picco might underestimate the finite-size correction, which finally caused the overestimate of $\eta$, and by studying the short-distance correction to the correlation function, they proposed that when properly considering the finite-size correction term, $\eta$ would revert to $0.25$ at $\sigma=1.75$, consistent with Sak’s criterion. Inspired by Ref. [9], we also check our estimate of $\eta$ at $\sigma=1.75$ following the same method by studying the correlation function. The analysis is conducted in the next subsection, and our estimate of $\eta$ is found to be reliable. Panel (b) in Fig. 17 demonstrates our numerical estimates of $y_{t}$, as well as the $\epsilon$ expansion, which results up to the second order [2]. As $\sigma$ approaches $2$, the numerical estimates of $y_{t}$ gradually deviate from the results of the $\epsilon$-expansion.
@@ VI.3 Subleading Magnetic Exponents
Strong correction to correlation functions $g(r)$ has been observed in the 2D LR Ising model, and the neglect of such corrections could lead to an overestimate of $\eta$ [9]. In this part, following the same routine in Ref. [9], we study the correlation function at the critical point for the 2D LR XY model. In particular, we focus on the case of $\sigma=1.75$ where $\eta$ is estimated to be $0.324(7)$, and similar procedures can apply to other $\sigma$. Our estimated value of $\eta$ is taken to be $0.32$, with the corresponding correction exponent $\omega$ in Eq. (41) being $0.45$. By carefully fitting the correlation function, we identify a subleading power-law correction likely arising from the subleading magnetic exponent $y_{h,2}$ of the transition [37]. The fitting results are compatible with our estimates of $\eta$, thus validating our previous analysis.
@@
where the summation only goes through $N_{s}$ randomly chosen sites instead of the entirety, and $N_{s}=100$ in our practice. In this way, the complexity of this measurement remains $O(N)$ instead of $O(N^{2})$, which allows us to measure the system with a larger size and simulate more samples. The cost is the slightly higher uncertainty in the measurement, but this can be compensated by involving more samples.
The left panel in Fig. 18 demonstrates the correlation function in various $L$ where the double-log scale is used to observe the power-law behavior. The inset shows the curves of the correlation function with the x-axis being distance $r$. It can be observed that curves flatten in their tails near $L/2$ due to the finite-size effect. To diminish the effect and for a better observation of the power-law behavior, we follow the treatment of Ref. [9] to choose the x-axis as $\sin(\pi r/L)L/\pi$ in the main figure. It does not affect the correlation function in short distances since $\sin(\pi r/L)L/\pi\approx(\pi r/L)L/\pi=r$. Moreover, as can be seen in the main figure, it drastically reduces the boundary effect. In both the inset and the main figure, $g(r)$ exhibits a fast decay in the short distance and then decays slowly as $r$ increases, which suggests an additional correction to the correlation function. We then perform a detailed fit to the correlation function of the largest system size $L=4096$ with $\eta$ fixed at $0.32$:
@@
The additional power-law correction to $g(r)$ should originate from a subleading relevant magnetic scaling field. In the context of RG, the free-energy density depends on not only the leading thermal and magnetic scaling fields $t_{1}\equiv t$, $h_{1}\equiv h$ but rather a full set of scaling fields. Each scaling field is associated with a particular eigenvalue of the RG transformation. In particular, the subleading thermal and magnetic scaling fields $t_{2}$ and $h_{2}$ correspond to the subleading eigenvalues in the even sector and odd sector, respectively. The contributions of these subleading exponents are very difficult to observe explicitly from numerical simulations because they either decay very fast or have very small, even zero amplitude. For instance, for the NN Ising model, one has $y_{t,2}=-4/3$ and $y_{h,2}=13/24\approx 0.542$ [54]. The subleading thermal field $t_{2}$ is irrelevant, whose contribution to FSS decays quickly. As for the subleading magnetic field $h_{2}$, despite being relevant, it is commonly believed to be redundant [55, 56]. To our knowledge, direct numerical evidence for $y_{h,2}$ remains obscure.
@@
We are grateful to observe the contribution of the subleading magnetic field at criticality in the LR XY model, which is difficult to identify numerically in previous studies of SR cases. In our opinion, this strong correction to $g(r)$ is another piece of evidence that the system is not in the SR universality class at $\sigma=1.75$. Note that the double power correlation function observed in the LR Ising model could also be attributed to the contribution of $y_{h,2}$, suggesting a different universality class than the SR one at $\sigma=1.75$.
@@ VII Conclusion
In this study, we have extended our previous work [19] to provide a comprehensive understanding of the two-dimensional long-range (LR) XY model with algebraically decaying interactions. We conduct large-scale Monte Carlo simulations to study the low-$T$, high-$T$, and critical properties of the model at various $\sigma$. We present strong evidence that the threshold between long-range and short-range universality lies at $\sigma_{*}=2$, which contrasts Sak’s scenario [57] and previous theoretical predictions [17, 18].
We show that for $\sigma\leq 2$, the system exhibits long-range order and Goldstone mode excitations in the low-$T$ phase, while for $\sigma>2$, the system enters a QLRO phase via a BKT transition, belonging to the SR universality. Furthermore, we reveal the distinct growth behaviors of the correlation length $\xi$ in the high-$T$ phase as the system approaches $T_{c}$. For $\sigma\leq 2$, the correlation length follows a power-law divergence, indicative of a second-order transition; for $\sigma>2$, the typical exponential growth of the BKT transition is observed. Finally, we determine the critical points and critical exponents of the model in the regime $1<\sigma\leq 2$. The resulting phase diagram highlights the different universality classes for $\sigma\leq 2$ and $\sigma>2$, providing deeper insights into the phase transitions in long-range interacting systems.
The findings presented in this work significantly enhance our understanding of the crossover between short-range and long-range interactions in the two-dimensional XY model. This crossover is crucial for elucidating the impact of long-range interactions in experimental systems, such as Rydberg atom arrays, where tunable interactions are prevalent. Moreover, our study provides valuable insights into how the nature of these interactions influences phase behavior and critical phenomena, offering broader implications for the fields of statistical mechanics, condensed matter physics, and cold atom physics. This work also lays the groundwork for future research into systems with long-range interactions, like the long-ranged Ising model and Heisenberg model, as well as the self-avoiding walk.
Acknowledgements

### v2

@@ VI.2 The Fit of Critical Exponents
where $b\cdot L^{-\omega}$ refers to the finite-size correction and $c$ derives from the analytic part of freedom energy. During the fitting process, various values of $\omega$ are tested. It is observed that when $\omega$ is excessively large or small, the data cannot be adequately fitted. However, when $\omega$ falls within a certain range, an appropriate fit yields reliable estimates for exponents. The details and results of the fitting are shown in Tables 6, 7, and 5, respectively. An interesting point is that our fitted $\eta$ values are consistent with previous results for LR Ising model [6, 9]. Figure 17 and 18 visually show the reliability of the fitting.
Table 5: Estimates of two exponents $y_{t}$ and $\eta$ for nonclassical regime. The results are obtained from the fitting: Eq. (43) and (44). Some of the estimated values of $\eta$ for LR Ising model by Luijten [6] and Picco [9] are also included for comparison.
$\sigma$       $y_{t}$       $\eta$ $\eta$ (Luijten) $\eta$ (Picco)
@@
2 0.37(4) 0.260(5) 0.266(16) 0.262(4)
To be specific, when taking the estimated value of $\eta$ and $y_{t}$, according to the fitting, $\chi L^{\eta-2}-bL^{-\omega}$ and $KL^{-y_{t}}-bL^{-\omega}$ should converge to a non-zero value as $L$ increases; otherwise they’ll either diverge or reduce to 0. It can be seen that in both figures, the curves corresponding to the estimated values of exponents converge to a plateau, indicating the reliability of the fitting results. Figure 19 presents our numerical estimates of critical exponents at different $\sigma$. Panel (a) displays our estimates for $\eta$ across different $\sigma$ values, alongside the estimates by Picco [9] and Luijten [6].
Figure 19: The value of critical exponent at various $\sigma$. (a) The red dots denote the estimates of $\eta$ in this work. Picco and Luijten’s results are also contained [9, 6], respectively denoted by blue and green dots. Although the model they studied is the Ising model, our estimated values of $\eta$ are consistent with their results. The dotted lines correspond to Sak’s prediction: $\eta=\rm{max}(2-\sigma,\frac{1}{4})$. Obviously, our results deviate from Sak’s prediction, suggesting that for $21.75<\sigma<$ 2, the short-range behavior is not recovered. (b) The dotted line represents the results from second-order $\epsilon$ expansion [2] (note that $y_{t}=1/\nu$). The red dots denote the estimates of $y_{t}$ in this work. Note that for $\sigma>2$, the LR system undergoes a BKT phase transition instead of a second-order one, with corresponding $y_{t}=0$. Therefore, there will be a jump for the value of $y_{t}$ at $\sigma=2$. This jump is reasonable considering the distinct different type of phase transition for $\sigma=2$ and $\sigma>2$.
@@
Despite the difference in models (Luijten and Picco studied the 2D LR Ising model, whereas we are studying the 2D LR XY model), our results are consistent with Picco’s, which does not agree with Sak’s prediction. In our results, $\eta$ interpolates smoothly between $\eta=2-\sigma$ for $\sigma$ near 1 and $\eta=\eta_{\text{SR}}=\frac{1}{4}$ for $\sigma>2$. For $\sigma<2$, $\eta$ remains significantly higher than $\frac{1}{4}$, indicating the phase transition doesn’t belong to the short-range universality class. For $\sigma=2$, the estimated value of $\eta$ still slightly exceeds $\frac{1}{4}$, which could be attributed to some logarithmic corrections in the behavior of $\chi$ at the marginal point $\sigma=2$. Panel (b) in Fig. 19 demonstrates our numerical estimates of $y_{t}$, as well as the $\epsilon$ expansion results up to the second order [2]. It can be seen that our numerical data starts to deviate from the $\epsilon$ expansion result from $\sigma\approx 1.5$. This should also apply to the value of $\eta$ according to the panel (a). Therefore, the $\epsilon$ expansion result near $\sigma=1$, $\eta=2-\sigma$ should not always hold along the whole nonclassical regime, especially for $\sigma\geq 1.5$ cases. However, Sak criterion simply adopt the relation for $\sigma\leq 1.75$, which we considered unreliable.
@@ VI.3 Subleading Magnetic Exponents
As mentioned before, Ref. [7] suggests that Picco might underestimate the finite-size correction, which finally caused the overestimate of $\eta$. By studying the short-distance correction to the correlation function, they proposed that when properly considering the finite-size correction term, $\eta$ would revert to $0.25$ at $\sigma=1.75$, consistent with Sak’s criterion. In this subsection, following the same routine in Ref. [7], we also study the correlation function at the critical point for the 2D LR XY model. In particular, we focus on the case of $\sigma=1.75$ where $\eta$ is estimated to be $0.324(7)$, and similar procedures can apply to other $\sigma$. Our estimated value of $\eta$ is taken to be $0.32$, with the corresponding correction exponent $\omega$ in Eq. (43) being $0.45$. By carefully fitting the correlation function, we identify a subleading power-law correction likely arising from the subleading magnetic exponent $y_{h,2}$ of the transition [39]. The fitting results are compatible with our estimates of $\eta$, thus validating our previous analysis.
@@
where the summation only goes through $N_{s}$ randomly chosen sites instead of the entirety, and $N_{s}=100$ in our practice. In this way, the complexity of this measurement remains $\mathcal{O}(N)$ instead of $\mathcal{O}(N^{2})$, which allows us to measure the system with a larger size and simulate more samples. The cost is the slightly higher uncertainty in the measurement, but this can be compensated by involving more samples.
Figure 22: The plot of $C_{\text{NN}}$ versus $T$ is shown for (a) $\sigma=1.25$, (b) $\sigma=1.75$, (c) $\sigma=2$, and (d) $\sigma=3$. The dashed vertical lines indicate the transition points for various $\sigma$ values. For $\sigma\leq 2$, finite yet sharp peaks are observed at the critical points, consistent with our prediction of second-order phase transitions. In contrast, the smooth and broad peak for $\sigma=3$ is the typical behavior of a BKT transition.
The inset of the left panel in Fig. 20 demonstrates the correlation function with various system size $L$ where the double-log scale is used to observe the power-law behavior. The flattening tails of the correlation functions are due to the finite-size effect. To diminish such effect for a better observation of the power-law behavior, we follow the treatment of Ref. [7] to choose the x-axis as $\sin(\pi r/L)L/\pi$ in the main figure. It does not affect the correlation function in short distances since $\sin(\pi r/L)L/\pi\approx(\pi r/L)L/\pi=r$. Moreover, as can be seen in the main figure, it drastically reduces the boundary effect. In the main figure, $g(r)$ exhibits a fast decay in the short distance and then decays slowly as $r$ increases, which suggests an additional correction to the correlation function. We then perform a detailed fit to the correlation function of the largest system size $L=4096$ with $\eta$ fixed at $0.32$:
@@
The additional power-law correction to $g(r)$ should originate from a subleading relevant magnetic scaling field. In the context of RG, the free-energy density depends on not only the leading thermal and magnetic scaling fields $t_{1}\equiv t$, $h_{1}\equiv h$ but rather a full set of scaling fields. Each scaling field is associated with a particular eigenvalue of the RG transformation. In particular, the subleading thermal and magnetic scaling fields $t_{2}$ and $h_{2}$ correspond to the subleading eigenvalues in the even sector and odd sector, respectively. The contributions of these subleading exponents are very difficult to observe explicitly from numerical simulations because they either decay very fast or have very small, even zero amplitude. For instance, for the NN Ising model, one has $y_{t,2}=-4/3$ and $y_{h,2}=13/24\approx 0.542$ [61]. The subleading thermal field $t_{2}$ is irrelevant, whose contribution to FSS decays quickly. As for the subleading magnetic field $h_{2}$, despite being relevant, it is commonly believed to be redundant [62, 63]. Recently, the exponent $y_{h,2}$ has been observed in the Fortuin-Kasteleyn geometric representation of the Ising model [64].
@@
We are glad to observe the contribution of the subleading magnetic field at criticality in the LR XY model, which is difficult to identify numerically in previous studies of SR cases. In our opinion, this strong correction to $g(r)$ is another piece of evidence that the system is not in the SR universality class at $\sigma=1.75$, since the two cases do not have the same subleading magnetic exponent. Note that the double power correlation function observed in the LR Ising model could also be attributed to the contribution of $y_{h,2}$, suggesting a different universality class with the SR one at $\sigma=1.75$.
VI.4 The properties of the specific-heat-like quantity
As a supplemental indication for the crossover from SR universality to LR universality, the specific heat $C$ exhibits somewhat different behaviors in the two regimes. In the SR regime, the specific heat $C$ exhibits a smooth and broad peak, which is a hallmark of the BKT transition. Conversely, as the system approaches a second-order phase transition, the singular part of $C$ scales $\sim\left|t\right|^{-\alpha}$, where $t$ is the reduced temperature and $\alpha$ is the specific-heat critical exponent. When $\alpha>0$, the specific heat diverges at the critical temperature $T_{c}$. However, for $\alpha<0$, $C$ remains finite at $T_{c}$ despite the presence of critical singularities. In such cases, the specific heat typically develops a sharp peak or even a nonanalytic kink at the transition point, which can serve as a hint of a second-order phase transition.
Directly measuring the full energy density and specific heat for long-range interacting systems is computationally expensive. Hence, we compute the energy of nearest neighbors for simplicity, defined as
$\varepsilon=L^{-2}\sum_{\langle ij\rangle}\bm{S}_{i}\cdot\bm{S}_{j},$ (47)
where the summation is over all NN pairs. This energy-like quantity shares the same critical scaling form of the energy density $\langle E\rangle$. Thus, it is natural to define a specific-heat-like quantity that possesses the identical scaling form of the specific heat,
$C_{\text{NN}}=\beta^{2}L^{2}\left(\langle\varepsilon^{2}\rangle-\langle\varepsilon\rangle^{2}\right),$ (48)
In Fig. 22, we plot $C_{\text{NN}}$ versus $T$ for $\sigma=1.25$, $1.75$, $2$, and $3$. For $\sigma=3$, $C_{\text{NN}}$ exhibits a smooth and broad peak slightly above the transition temperature $T_{c}$, which is a typical feature of the BKT transition. In contrast, for $\sigma=1.25$, $1.75$, $2$, finite yet sharp peaks are clearly observed near $T_{c}$. As the $\sigma$ decreases, the peak becomes sharper and more asymmetric, indicating nonanalytic behavior in the free energy. This observation is consistent with our analysis of critical exponents. For a system below its upper critical dimension, the specific-heat critical exponent $\alpha$ is related to the correlation length exponent $\nu$ via the hyperscaling relation $\alpha=2-d\nu$. Table 5 shows that $y_{t}<1$ for $\sigma=1.25$, $1.75$, $2$, which implies $11/y_{t}=\nu>$ 1 and $\alpha<0$. Therefore, at these $\sigma$ values, the specific heats develop cusps at the critical points.
Moreover, since much stronger evidence has already been provided from the low-T properties and the power-law divergence of $\xi$ as the critical point is approached from the high-T side, a comprehensive analysis of $C_{\text{NN}}$ is beyond the scope of this work. We note that, even though $C_{\text{NN}}$ alone does not provide definitive evidence for locating the crossover point $\sigma_{*}$, it offers valuable supplementary insight. The distinct behaviors observed in the $\sigma\leq 2$ and $\sigma>2$ regimes are consistent with our previous analysis and provide additional support for the existence of a crossover at $\sigma=2$.
@@ VII Conclusion
In this study, we provide a comprehensive understanding of the two-dimensional long-range (LR) XY model with algebraically decaying interactions. We perform large-scale Monte Carlo simulations to study the low-T, high-T, and critical properties of the model at various $\sigma$. Our results provide compelling evidence that the threshold between long-range and short-range universality lies at $\sigma_{*}=2$, in contrast to Sak’s scenario [65] and previous theoretical predictions [19, 20]. The main findings are summarized below:
  1. 1.
    For $\sigma\leq 2$, the system is shown to exhibit long-range order and Goldstone mode excitations in the low-T phase, while for $\sigma>2$, the system exhibits quasi-long-range order as in the NN case.
  2. 2.
    We reveal the distinct growth behavior of the correlation length $\xi$ in the high-T phase as the system approaches $T_{c}$. For $\sigma\leq 2$, the correlation length follows a power-law divergence, the signature of a second-order transition; for $\sigma>2$, the typical exponential growth of the BKT transition is observed.
  3. 3.
    The critical points and critical exponents of the model in the regime $1<\sigma\leq 2$ are determined. The estimates of $y_{t}$ and $\eta$ start to deviate from the $\epsilon$-expansion result from approximately $\sigma=1.50$, and the deviation becomes bigger as $\sigma$ further increases. This strongly suggests that the $\epsilon$-expansion results, like $\eta=2-\sigma$, are perturbative and should not be regarded as quantitatively true, particularly when $\sigma$ approaches 2. Moreover, our estimated values of critical exponents show distinct difference for $\sigma\leq 2$ and $\sigma>2$.
[…]
Acknowledgements.
@@
Appendix A: Extrapolation Detail
Table 8: Specific values of parameters obtained from the fit of Eq. (39) for $n=9$ and different $(L_{\mathrm{min1}},L_{\mathrm{min2}})$.
$L_{\text{min}}$ $a_{1}(\times 10^{0})$ $a_{2}(\times 10^{0})$ $a_{3}(\times 10^{2})$ $a_{4}(\times 10^{3})$ $a_{5}(\times 10^{3})$ $b$ $\omega$ $\chi^{2}$/DF
(128, 256) 2.22(18) -8.2(49) 2.83(51) -1.30(23) 1.63(36) 1.44(13) 0.538(19) 191.1/93
(128, 512) 2.24(18) -12.2(48) 3.31(49) -1.51(21) 1.96(33) 1.37(10) 0.514(16) 91.6/71
(256, 256) 2.06(19) -2.7(51) 2.15(55) -0.97(25) 1.10(39) 1.12(15) 0.494(25) 148.3/80
(256, 512) 2.14(17) -7.4(47) 2.70(48) -1.22(21) 1.49(33) 1.30(16) 0.505(23) 55.9/58
(512, 256) 2.07(20) -3.4(58) 2.20(63) -0.98(29) 1.09(47) 1.01(15) 0.476(27) 141.6/73
(512, 512) 2.11(20) -6.4(55) 2.57(59) -1.16(27) 1.39(42) 1.23(24) 0.496(34) 54.9/51
Table 9: Specific values of parameters obtained from the fit of Eq. (39) for $n=9$ and different $(L_{\mathrm{min1}},L_{\mathrm{min2}})$.
$L_{\text{min}}$ $a_{1}(\times 10^{0})$ $a_{2}(\times 10^{1})$ $a_{3}(\times 10^{3})$ $a_{4}(\times 10^{4})$ $a_{5}(\times 10^{5})$ $a_{5}(\times 10^{5})$ $a_{5}(\times 10^{5})$ $\chi^{2}$/DF
(64, 256) 2.14(33) -5.3(20) 1.57(44) -1.79(47) 1.07(26) -3.17(71) 3.64(77) 113.2/50
(64, 512) 2.26(27) -6.1(16) 1.78(37) -2.03(40) 1.20(22) -3.49(60) 3.94(66) 68.1/46
(128, 256) 1.66(35) -2.3(22) 0.90(48) -1.07(52) 0.67(28) -2.11(78) 2.53(84) 79.3/41
(128, 512) 1.82(24) -3.5(15) 1.19(34) -1.40(36) 0.86(20) -2.62(55) 3.05(60) 33.0/37
(256, 256) 1.62(55) -1.7(32) 0.69(67) -0.80(69) 0.51(37) -1.64(99) 2.0(10) 63.6/29
(256, 512) 1.85(33) -3.3(19) 1.06(41) -1.21(42) 0.74(23) -2.25(61) 2.64(64) 19.1/25
As mentioned in the main text, using the FSS ansatz (38), we extrapolate the correlation length $\xi$ to its thermodynamic-limit value. For $\sigma=2$ and $3$, the ansatz (38) is fitted to Eq. (39) and (40) respectively. To further account for the finite-size correction, we gradually increase the minimum system size $L_{\text{min}}$ included in the fitting. It is found that for both cases, finite-size corrections are more pronounced in the range $\xi/L\in(0,0.35)\cup(0.65,1)$ than in other intervals. Therefore, two different minimum system sizes $L_{\text{min1}}$ and $L_{\text{min2}}$ are selected for two separate intervals, namely the range $\xi/L\in(0.35,0.65)$ and $\xi/L\in(0,0.35)\cup(0.65,1)$ respectively, with $L_{\text{min1}}\leq L_{\text{min2}}$. Fitting results with different $(L_{\text{min1}},L_{\text{min2}})$ for $\sigma=2$ and $3$ are presented in Table 8 and 9 respectively. The extrapolation results in Fig. 14 use the optimal fit with $(L_{\text{min1}}=256,L_{\text{min2}}=512)$ for $\sigma=2$ and $(L_{\text{min1}}=128,L_{\text{min2}}=512)$ for $\sigma=3$. Further extrapolation details can be found in Ref. [44]